INAF queue observing, 2018 Mar 23-24

Observer: R. Carini, A. Rossi
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: B Rothberg on call
Instrument: MODS1+MODS2, LBC

Summary

At the sunset the sky is whispy clouds and high wind but still within limit
instruments: LBCs, MODS1 and LUCIs work well, MODS2 is working with blue only, LUCI1 and LUCI2 are ok
Night duration (12\xB0 el twilight) 2:20am- 12:35pm UT = 10 h 15m

Observing time: 2h 15m (22%)

Weather time loss: 0h (78%)

technical time loss 0 (0%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibrations dark LUCI1 + LUCI2 camera N3.75      
Calibrations flat LUCI1 + LUCI2 N3.75, FeII, J Ks      
Calibrations flat LUCI1 + LUCI2 N30 K, Ks      
Calibration arc MODS2 blue Xe+Kr      
Antoniucci DrTau LUCI1+2 longslit zJ HK 0.75''   luci1.0027-0034==, ==luci2.0027-0034 completed Abbiamo rispettato i tempi dati nello script, non nel proposal
Antoniucci DrTau LUCI1+2 imaging FeII 0.1/0.1 luci1.0027-0034==, ==luci2.0027-0034 completed Abbiamo rispettato i tempi dati nello script, non nel proposal

Details

LUCI
  • dark luci1.201803234.0006-0010; luci2.20180324.0006-0010 dit= 5, ndit=6, int, lir. agll354
  • dark luci1.201803234.0011-0015; luci2.20180324.0011-0015 dit= 300, ndit=1, norm, mer. agll354
  • flats luci1.201803234.0041-0050; luci2.20180324.0041-0050 flats FeII for Antoniucci imaging, no good lights one in chamber
  • flats luci1.201803234.0051-0060; luci2.20180324.0051-0060 flats J for imaging, no good lights one in chamber
  • flats luci1.201803234.0061-0070; luci2.20180324.0061-0070 flats Ks for imaging, no good lights one in chamber
  • flats luci1.201803234.0082-0086; luci2.20180324.0086-0090 flats Ks for N30 camera
  • flats luci1.201803234.0092-0096; luci2.20180324.0096-0100 flats K for N30 camera
  • flats luci1.201803234.0103-0107; luci2.20180324.0107-0111 flats FeII for Antoniucci imaging, good
  • flats luci1.201803234.0113-0117; luci2.20180324.0117-0121 flats J for imaging, good
  • flats luci1.201803234.0123-0127; luci2.20180324.0127-0131 flats Ks for imaging,good

MODS
  • arc Xe+Kr Lamps in MODS2 blue mods2.20180324.0001-0002

02:00 Since variable weather conditions and the presence of clouds we start the night with a back-up program.
  • say to TO to configure LUCI 1 and LUCI2

02:14 start with telluric HIP31128
  • acq luci1.0016-18, luci2.0016-18
  • seeing 1" at 02:23
  • sci luci1.0019-0022, luci2.0019-0022

02:29 stop for collimating guiding again

02:51 Dome closed for wind

03:20 open the dome
  • 03:45 begin the science, Antoniucci's program, target Dr Tau, seeing about 1.3''
  • acq luci1.0023-0026, luci2.0023-0026
  • sci luci1.0027-0034, luci2.0027-0034

04:10 we slew to MODS1 + LBC red because the seeing is very good, even if there are cirrus

:04:44 Gilli program, mask 31
  • acq mods1r.0002-04
  • seeing 1" at 5:00 acquisition and allignement done w MODS1
  • 05:08 dohybrid failed, try dofpia, \first, \Redonly
  • 05:12 start MODS1 science
  • sci mods1r.0005,mods1b.0002
  • seeing 1" at 5:15 0.2-0.3 extinction on DIMM
  • 05:21 dofpia did not converge, clear adsec, try again

05:27 guide lost due the clouds, stop integration and dofpia.
  • wind gusts rises the limit of about 22 m/s, dome closed
  • the images saved are of 516 sec in blue and 506 sec in red

06:10 strong wind and clouds now, we stay closed

11:00 sky is clear now but wind too strong

11:50 TO opens the dome, the sky is clearer and the wind is decreased.
  • the target Caccianiga H is in front of the wind, we can not observe it. The wind is again too near to the limit, we close!!!!

-- %USERSIG{AndreaRossi - 2018-03-24}%

Comments

 
Topic revision: r13 - 26 Mar 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback